Bartolo Delivery has two divisions, air express and ground service, that share the common costs of the company’s communications network, which are $8,900,000 a year. You have the following information about the two divisions and the common communications network:

Calls (thousands) Time on Network (hours)
Air express 588,000 337,500
Ground service 252,000 1,012,500

Required:

Determine the cost allocation if $6.23 million of the communications network costs are fixed and allocated on the basis of time on network, and the remaining costs, which are variable, are allocated on the basis of the number of calls. (Enter your answers in whole dollars. Do not round intermediate calculations.)
Answer #1

Allocation of fixed cost s 6.23 millions to

Air express = 6230000* (337500/(337500+1012500) = 1557500

Ground service = 6230000 * ( 1012500/(337500+1012500)) = 4672500

Variable cost

= 8900000 - 6230000 ( total cost - fixed cost)

= 2670000

Allocation of variable costs to

Air express = 2670000 * (588000/(588000+252000)) = 1869000

Ground service = 2670000 * ( 252000/(588000+252000)) = 801000